OpenStudy (anonymous):

Given the length of a line (d) and an angle (θ), sin and cos can be used to extract the x and y components of the line. \[d_{component}=d*\sin(\theta) \quad \bigg| \quad d_{component}=d*\cos(\theta)\] Which function gives the x component and which gives the y component? Why? (It's been a while since I've used these.)

OpenStudy (anonymous):

I am more used to seeing this in the context of vectors. If we have a triangle with your d as the hypotenuse and a forming an angle theta with d in the x direction and b the other side in the y-direction, then Sin theta = b/d and Cos theta = a/d.

OpenStudy (anonymous):

Thank you much.